# Principled RLHF from Heterogeneous Feedback via Personalization and Preference Aggregation

### v1

@@ Acknowledgement
The authors would like to thank Dingwen Kong for the discussions, and thank Jisu Jang for providing helpful feedback for Figure 1.
@@
“Principled RLHF from Heterogeneous Feedback via
@@
Appendix A Table of Notation
@@
Appendix B Deferred Definition
B.1 Bracketing Number.
@@ Appendix B Deferred Definition
B.2 Concentrability Coefficient for a Reward Scalar Class
@@
Appendix C Deferred Proofs in Section 3
C.1 Proof of Theorem 3.1 and Corollary 3.2
See 3.1 See 3.2 Before having a proof of Theorem 3.1 and Corollary 3.2, we provide general two properties of MLE estimates, which is a slightly modified version of (Zhan et al., 2023) and (Liu et al., 2022).
@@ Appendix C Deferred Proofs in Section 3
Note that $\bm{r}^{\star}$ do not need to be in $\mathcal{G}_{\bm{r}}^{\prime}$ for the above lemmas. Lemma 2 states that the log-likelihood $\log P_{\bm{r}}$ for a preference dataset generated by the reward model $\bm{r}^{\star}$ cannot exceed the log-likelihood $\log P_{\bm{r}^{\star}}$ for a preference dataset generated by the reward model $\bm{r}^{\star}$, with a gap related to the bracket number of $\mathcal{G}_{\bm{r}}$. Lemma 4 states that the $\ell_{1}$ distance between likelihood function $P_{\bm{r}^{\star}}$ and $P_{\bm{r}}$ for all $\bm{r}\in\mathcal{G}_{\bm{r}}^{\prime}$ can be bounded with the difference between log-likelihood $\log P_{\bm{r}^{\star}}$ and $\log P_{\bm{r}}$ for a preference dataset generated by the reward model $\bm{r}^{\star}$ with a gap related to the bracket number of $\mathcal{G}_{\bm{r}}^{\prime}$.
Proof of Theorem 3.1 and Corollary 3.2.
@@
which indicates that $\bm{r}^{\star}(=\bm{r}_{\omega^{\star},\boldsymbol{\theta}^{\star}})\in\mathcal{R}(\widehat{\mathcal{D}})$. Moreover, by the definition of Equation 3.1, if $\bm{r}_{\omega,\boldsymbol{\theta}},\bm{r}_{\omega^{\prime},\boldsymbol{\theta}^{\prime}}\in\mathcal{R}(\widehat{\mathcal{D}})$,
@@
$\frac{1}{N}\sum_{i\in[N]}\mathbb{E}_{\mu_{0},\mu_{1}}\left[\left|(r_{\omega,\theta_{i}}(\tau_{i,0})-r_{\omega,\theta_{i}}(\tau_{i,1}))-(r_{i}^{\star}(\tau_{i,0})-r_{i}^{\star}(\tau_{i,1}))\right|^{2}\right]$ (C.1)
@@
$J(\pi_{i,\text{tar}};r^{\star}_{i})-J(\widehat{\pi}_{i};r^{\star}_{i})$ (C.2)
@@
Here, $(i)$ holds since $\widehat{\pi}_{j}$ is a distributional robust policy for $\mathcal{R}(\widehat{\mathcal{D}})$ (Equation 3.1) and $(ii)$ holds due to the definition of $r_{\widehat{\pi}_{i}}^{i,\text{inf}}$. Therefore, if we sum Equation C.2 over $i\in[N]$, we have
@@
which proves Theorem 3.1. Moreover, we have
@@
which proves Corollary 3.2. ∎
C.2 Discussion on 3.3
C.2.1 Comparing with (Lu et al., 2021, Assumption 6.4)
Assumption C.1 ((Lu et al. (2021), Assumption 6.4)).
@@
3.3 bears similarity to C.1; however, the latter is notably more stringent. For instance, consider the case where $v=v^{\prime}=e_{1}$ without loss of generality. If it holds that $\mathbb{E}\|\psi_{1}(x)-\psi^{\prime}_{1}(x)\|^{2}\leq\epsilon$, then it implies $\psi\sim P\psi^{\prime}$. In this context, $\psi_{1}$ and $\psi^{\prime}_{1}$ represent the first coordinates of $\psi$ and $\psi^{\prime}$, respectively. The assumption that similarity in the first coordinate necessitates equivalence of the entire representations ($\psi\sim P\psi^{\prime}$) is a strong assumption.
C.2.2 Case Study (Linear Representation): $\psi_{\omega}(x)=\omega x$ and $\omega$ is an Orthonormal Matrix
@@
$\frac{1}{T}\sum_{i\in[T]}\mathbb{E}_{\tau\sim\mu}\|\psi_{\omega}(\phi(\tau))^{\top}v_{i}-\psi_{\omega^{\prime}}(\phi(\tau))^{\top}v_{i}^{\prime}\|^{2}\leq\epsilon$ (C.3)
@@
By Equation C.1, if we use 3.3 with $\Theta^{\star}/B$, we can find an orthonormal matrix $P_{\omega}$ such that
@@
Suppose Assumptions 3.1, 3.2 and 3.3 hold. For any $\delta\in(0,1]$ and $\lambda>0$, with probability at least $1-\delta$, $\bm{r}^{\star}\in\mathcal{R}^{\prime}(\widehat{\mathcal{D}})$, i.e., the underlying reward functions are an element of Equation 3.2.
@@
$\left[\|\psi^{\star}(\phi(\tau_{0}))-\psi^{\star}(\phi(\tau_{1}))-P_{\widehat{\omega}}(\psi_{\widehat{\omega}}(\phi(\tau_{0}))-\psi_{\widehat{\omega}}(\phi(\tau_{1})))\|^{2}\right]\leq k\frac{c_{\text{rep}}\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{{NN_{p}B^{2}}}.$ (C.4)
We only consider the event that Equation C.4 holds. We will use this $P_{\widehat{\omega}}$ for the proof of Theorem 3.4. We will approach similarly with the proof of (Zhu et al., 2023). Consider the following optimization problem:
@@
$\|\nabla f(P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star})\|_{(\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I)^{-1}}\|\widehat{\theta}_{i}-P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}\|_{\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I}\geq\langle\nabla f(P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}),\widehat{\theta}_{i}-P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}\rangle\geq\frac{\eta}{2}\|\widehat{\theta}_{i}-P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}\|^{2}_{\widehat{\Sigma}_{\psi_{\widehat{\omega}}}}.$ (C.5)
@@
$\|\frac{1}{N_{p}}\psi_{\widehat{\omega}}(\widehat{\mathcal{D}}_{i})^{\intercal}V_{i}\|_{(\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I)^{-1}}\leq\xi C_{4}\sqrt{\frac{k+\log(N/\delta)}{N_{p}}}$ (C.6)
@@
${\|\frac{1}{N_{p}}\psi_{\widehat{\omega}}(\widehat{\mathcal{D}}_{i})^{\intercal}(V_{i}^{\prime}-V_{i})\|_{(\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I)^{-1}}}\leq\frac{\xi C_{5}}{\sqrt{N_{p}}}\sqrt{k\frac{\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{{NN_{p}}}}$ (C.7)
@@
Combining Equation C.6 and Equation C.7, we have
@@
for a constant $C_{6}>0$ with probability at least $1-\delta/N$ and Equation C.5 provides
@@
Suppose Assumptions 3.1, 3.2, 3.3, and 3.4 hold. For any $\delta\in(0,1]$, with probability at least $1-\delta$, for any $\bm{r}_{\omega,\boldsymbol{\theta}}\in\mathcal{R}^{\prime}(\widehat{\mathcal{D}})$,
@@
For any $\tau_{0},\tau_{1}$, by 3.4, with large $N_{p}\geq N_{\text{unif}}(\Psi,\mu_{0},\mu_{1},\delta)$, we have the analog of Equation C.1:
@@
where $c>0$ is a constant, which is similar to the proof of Theorem 3.1. ∎
@@
We follow the construction in Theorem 3.10 of Zhu et al. (2023).
@@
According to Zhu et al. (2023), $\left\|\Sigma_{\mathcal{D}}^{-1/2}\mathbb{E}_{s\sim\rho}\left[\psi_{\omega}(\phi(s,\pi^{\star}(s)))\right]\right\|_{2}\leq\Lambda$, where $\rho$ is the uniform distribution over ${\mathcal{S}}$. At the same time, for any $\theta_{w}$ we have $\left\|\theta_{w}\right\|_{2}\in\Theta_{B}$ when taking $B=1$, $d>6$ and $\Delta<1/(6d)$.
Next, we will show that $C_{\bm{r}}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\Lambda$. By definition, we have
@@
since $a^{\prime}\equiv a_{3}$ by definition of $\mu_{1}$ and $\psi_{\omega}(\phi(\cdot,a_{3})\equiv 0$. Then, by Section D.1. of Zhan et al. (2023), we have $C_{\bm{r}}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\left\|\Sigma_{\mathcal{D}}^{-1/2}\mathbb{E}_{s\sim\rho}\left[\psi_{\omega}(\phi(s,\pi^{\star}(s)))\right]\right\|_{2}\leq\Lambda$. Therefore, combined with Theorem 3.10 of Zhu et al. (2023), we finished the proof. ∎
Appendix D Proof of Section 4
@@
where $\xi:=\max_{x\in[-R_{\max},R_{\max}]}\left|\frac{\Phi^{\prime}(x)}{\Phi(x)}\right|$, which is also defined in Appendix C. ∎
See 4.1
@@
Here, we used $\mathcal{N}_{\mathcal{G}_{\psi_{\widehat{\omega}}}}(1/NN_{p})=\mathcal{N}_{\mathcal{G}_{\psi^{\star}}}(1/NN_{p})$. Now, we get the similar bound with Equation C.1:
@@
Appendix E Proof of Section 5
E.1 Six Pivotal Axioms for Reward Aggregation

### v2

@@ Acknowledgement
The authors would like to thank Jisu Jang for providing helpful feedback for Figure 1.
@@
“RLHF from Heterogeneous Feedback via
@@
Appendix A Societal Impact
Our work is mainly theoretical, and aimed at better understanding RLHF with heterogeneous feedback, with principles, algorithms, and analyses. As such, we do not anticipate any direct positive or negative societal impact from this research.
Appendix B Limitations
Our works provided overall theoretical analysis and experimental validation. However, due to the computational issue, we experimented on the 6B and 8B models, and also we did not calculate the penalty for the pessimism in our Algorithms.
Appendix C Table of Notation
@@
Appendix D Deferred Definition
Bracketing Number.
@@ Appendix D Deferred Definition
Concentrability Coefficient for a Reward Scalar Class
@@
Direct Preference Optimization (DPO) (Rafailov et al., 2024).
[…]
Appendix E Deferred Pseudocode of Algorithms
Algorithm 6 ClusterDPO: Learning $K$ clustered policies by DPO
  Randomly initialize $f^{0}(i)$ for $i\notin\{p_{1},\dots,p_{K}\}$
  for $t\in[T]$ do
     Randomly select $K$ human users $p_{1},\dots,p_{K}$.
     for $i\in[N]$ do
        if $i\notin\{p_{1},\dots,p_{K}\}$ then
           Define $f^{t}(i)\leftarrow f^{t-1}(i)$
        end if
     end for
     Assign $f^{t}(p_{k})$ for all $k\in[K]$ as
$f^{t}(p_{k})\leftarrow\underset{s\in[K]}{\arg\max}\sum_{j\in[N_{p}]}\log\sigma\left(\beta\log\frac{\pi_{(s)}^{t-1}(a_{p_{k},0}^{(j)}\mid s_{p_{k}}^{(j)})}{\pi^{\text{old}}(a_{p_{k},0}^{(j)}\mid s_{p_{k}}^{(j)})}-\beta\log\frac{\pi_{(s)}^{t-1}(a_{p_{k},1}^{(j)}\mid s_{p_{k}}^{(j)})}{\pi^{\text{old}}(a_{p_{k},1}^{(j)}\mid s_{p_{k}}^{(j)})}\right)$ (E.1)
and obtain $\pi_{(s)}^{t}$ for all $s\in[K]$.
  end for
  Assign $f^{T+1}(i)$ for all $i\in[N]$ as
  Output: $({\pi}_{(k)}^{T})_{k\in[K]}$ and $f^{T+1}$
Appendix F Deferred Proofs in Section 3.1
F.1 Expected Value Function Gap without Diversity Assumption
Firstly, we provide an algorithm for each reward function learning without Assumptions 2, 3, and 4.
Algorithm 7 Personalized RLHF via Representation Learning - without Diversity Assumption
  Estimate $\omega^{\star}$ and $\boldsymbol{\theta}^{\star}$ by
$(\widehat{\omega},\widehat{\boldsymbol{\theta}})\leftarrow\operatorname*{arg\,max}_{\omega\in\Omega,\|\theta_{i}\|_{2}\leq B\text{ for all }i\in[N]}\sum_{i\in[N]}\sum_{j\in[N_{p}]}\log P_{\omega,\theta_{i}}(o_{i}^{(j)}\mid\tau_{i,0}^{(j)},\tau_{i,1}^{(j)})$
  Construct a confidence set of the reward function by
[…]
$\widehat{\pi}_{i}\leftarrow\operatorname*{arg\,max}_{\pi\in\Pi}\min_{\bm{r}\in\mathcal{R}(\widehat{\mathcal{D}})}\left(J(\pi;r_{i})-\mathbb{E}_{\tau\sim\mu_{i,\text{ref}}}[r_{i}(\tau)]\right)$ (F.2)
  Output: $(\widehat{\omega},\widehat{\boldsymbol{\theta}},(\widehat{\pi}_{i})_{i\in[N]})$.
  • Confidence set (Equation F.1) for the MLE estimation as (Liu et al., 2022), which is also used in (Liu et al., 2023; Zhan et al., 2023; Wang et al., 2024; Zhan et al., 2022), with $\zeta=C_{1}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)$ for a constant $C_{1}>0$, which will be related to Theorem F.1. the definition of bracketing number ($\mathcal{N}_{\mathcal{G}_{\bm{r}}}$) is deferred to Appendix D.
We will provide the expected value function gap of the output of Algorithm 7 and the reference policy.
Theorem F.1.
(Total Expected Value Function Gap). Suppose 1 holds. For any $\delta\in(0,1]$, with probability at least $1-\delta$, the output $(\widehat{\pi}_{i})_{i\in[N]}$ of Algorithm 1 satisfies
Corollary F.1.
(Expected Value Function Gap). Suppose 1 holds. For any $\delta\in(0,1]$ and all $i\in[N]$, with probability at least $1-\delta$, the output $\widehat{\pi}_{i}$ of Algorithm 1 satisfies
where $c>0$ is a constant.
Note that the results above do not need any assumption on $(\theta_{i}^{\star})_{i\in[N]}$. Still, as $N_{p}\to\infty$, $\widehat{\pi}_{i}$ has comparable or better performance than the comparator policy $\pi_{i,\text{tar}}$, which approaches the optimal policy if $\pi_{i,\text{tar}}=\pi_{i}^{\star}$. We will leverage the proof of Theorem F.1 to prove Theorem 3.1. To be specific, we will improve the bound for Corollary F.1, as the gap of the expected value function does not decay with $N$, which is the number of human users. We defer the proofs of Theorem F.1 and Corollary F.1 to Section F.2.
F.2 Proof of Theorem F.1 and Corollary F.1
See F.1 See F.1 Before having a proof of Theorem F.1 and Corollary F.1, we provide two general properties of MLE estimates, which is a slightly modified version of (Zhan et al., 2023) and (Liu et al., 2022).
@@ F.1 Expected Value Function Gap without Diversity Assumption
Note that $\bm{r}^{\star}$ does not need to be in $\mathcal{G}_{\bm{r}}^{\prime}$ for the above lemmas. Lemma 2 states that the log-likelihood $\log P_{\bm{r}}$ for a preference dataset generated by the reward model $\bm{r}^{\star}$ cannot exceed the log-likelihood $\log P_{\bm{r}^{\star}}$ for a preference dataset generated by the reward model $\bm{r}^{\star}$, with a gap related to the bracket number of $\mathcal{G}_{\bm{r}}$. Lemma 4 states that the $\ell_{1}$ distance between likelihood function $P_{\bm{r}^{\star}}$ and $P_{\bm{r}}$ for all $\bm{r}\in\mathcal{G}_{\bm{r}}^{\prime}$ can be bounded with the difference between log-likelihood $\log P_{\bm{r}^{\star}}$ and $\log P_{\bm{r}}$ for a preference dataset generated by the reward model $\bm{r}^{\star}$ with a gap related to the bracket number of $\mathcal{G}_{\bm{r}}^{\prime}$.
Proof of Theorem F.1 and Corollary F.1.
@@
which indicates that $\bm{r}^{\star}(=\bm{r}_{\omega^{\star},\boldsymbol{\theta}^{\star}})\in\mathcal{R}(\widehat{\mathcal{D}})$. Moreover, by the definition of Equation F.1, if $\bm{r}_{\omega,\boldsymbol{\theta}},\bm{r}_{\omega^{\prime},\boldsymbol{\theta}^{\prime}}\in\mathcal{R}(\widehat{\mathcal{D}})$,
@@
$\frac{1}{N}\sum_{i\in[N]}\mathbb{E}_{\mu_{0},\mu_{1}}\left[\left|(r_{\omega,\theta_{i}}(\tau_{i,0})-r_{\omega,\theta_{i}}(\tau_{i,1}))-(r_{i}^{\star}(\tau_{i,0})-r_{i}^{\star}(\tau_{i,1}))\right|^{2}\right]$ (F.3)
@@
$J(\pi_{i,\text{tar}};r^{\star}_{i})-J(\widehat{\pi}_{i};r^{\star}_{i})$ (F.4)
@@
Here, $(i)$ holds since $\widehat{\pi}_{j}$ is a distributional robust policy for $\mathcal{R}(\widehat{\mathcal{D}})$ (Equation F.1) and $(ii)$ holds due to the definition of $r_{\widehat{\pi}_{i}}^{i,\text{inf}}$. Therefore, if we sum Equation F.4 over $i\in[N]$, we have
@@
which proves Theorem F.1. Moreover, we have
@@
which proves Corollary F.1. ∎
F.3 Discussion on 3
F.3.1 Comparing with (Lu et al., 2021, Assumption 6.4)
Assumption 5 ((Lu et al. (2021), Assumption 6.4)).
@@ F.3 Discussion on 3
3 bears similarity to 5; however, the latter is notably more stringent. For instance, consider the case where $v=v^{\prime}=e_{1}$ without loss of generality. If it holds that $\mathbb{E}\|\psi_{1}(x)-\psi^{\prime}_{1}(x)\|^{2}\leq\epsilon$, then it implies $\psi\sim P\psi^{\prime}$. In this context, $\psi_{1}$ and $\psi^{\prime}_{1}$ represent the first coordinates of $\psi$ and $\psi^{\prime}$, respectively. The assumption that similarity in the first coordinate necessitates equivalence of the entire representations ($\psi\sim P\psi^{\prime}$) is a strong assumption.
F.3.2 Case Study (Linear Representation): $\psi_{\omega}(x)=\omega x$ and $\omega$ is an Orthonormal Matrix
@@
$\frac{1}{T}\sum_{i\in[T]}\mathbb{E}_{\tau\sim\mu}\|\psi_{\omega}(\phi(\tau))^{\top}v_{i}-\psi_{\omega^{\prime}}(\phi(\tau))^{\top}v_{i}^{\prime}\|^{2}\leq\epsilon$ (F.5)
@@
By Equation F.3, if we use 3 with $\Theta^{\star}/B$, we can find an orthonormal matrix $P_{\omega}$ such that
@@
Suppose Assumptions 1, 2 and 3 hold. For any $\delta\in(0,1]$ and $\lambda>0$, with probability at least $1-\delta$, $\bm{r}^{\star}\in\mathcal{R}^{\prime}(\widehat{\mathcal{D}})$, i.e., the underlying reward functions are an element of Equation 3.1.
@@
$\left[\|\psi^{\star}(\phi(\tau_{0}))-\psi^{\star}(\phi(\tau_{1}))-P_{\widehat{\omega}}(\psi_{\widehat{\omega}}(\phi(\tau_{0}))-\psi_{\widehat{\omega}}(\phi(\tau_{1})))\|^{2}\right]\leq k\frac{c_{\text{rep}}\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{{NN_{p}B^{2}}}.$ (F.6)
We only consider the event that Equation F.6 holds. We will use this $P_{\widehat{\omega}}$ for the proof of Theorem 3.1. We will approach similarly with the proof of (Zhu et al., 2023). Consider the following optimization problem:
@@
$\|\nabla f(P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star})\|_{(\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I)^{-1}}\|\widehat{\theta}_{i}-P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}\|_{\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I}\geq\langle\nabla f(P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}),\widehat{\theta}_{i}-P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}\rangle\geq\frac{\eta}{2}\|\widehat{\theta}_{i}-P_{\widehat{\omega}}^{\intercal}\theta_{i}^{\star}\|^{2}_{\widehat{\Sigma}_{\psi_{\widehat{\omega}}}}.$ (F.7)
@@
$\|\frac{1}{N_{p}}\psi_{\widehat{\omega}}(\widehat{\mathcal{D}}_{i})^{\intercal}V_{i}\|_{(\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I)^{-1}}\leq\xi C_{4}\sqrt{\frac{k+\log(N/\delta)}{N_{p}}}$ (F.8)
@@
${\|\frac{1}{N_{p}}\psi_{\widehat{\omega}}(\widehat{\mathcal{D}}_{i})^{\intercal}(V_{i}^{\prime}-V_{i})\|_{(\widehat{\Sigma}_{\psi_{\widehat{\omega}}}+\lambda I)^{-1}}}\leq\frac{\xi C_{5}}{\sqrt{N_{p}}}\sqrt{k\frac{\kappa^{2}\log(\mathcal{N}_{\mathcal{G}_{\bm{r}}}(1/(NN_{p}))/\delta)}{{NN_{p}}}}$ (F.9)
@@
Combining Equation F.8 and Equation F.9, we have
@@
for a constant $C_{6}>0$ with probability at least $1-\delta/N$ and Equation F.7 provides
@@
Suppose Assumptions 1, 2, 3, and 4 hold. For any $\delta\in(0,1]$, with probability at least $1-\delta$, for any $\bm{r}_{\omega,\boldsymbol{\theta}}\in\mathcal{R}^{\prime}(\widehat{\mathcal{D}})$,
@@
For any $\tau_{0},\tau_{1}$, by 4, with large $N_{p}\geq N_{\text{unif}}(\Psi,\mu_{0},\mu_{1},\delta)$, we have the analog of Equation F.3:
@@
where $c>0$ is a constant, which is similar to the proof of Theorem F.1. ∎
@@
We follow the construction in Theorem 3.10 of (Zhu et al., 2023).
@@
According to (Zhu et al., 2023), $\left\|\Sigma_{\mathcal{D}}^{-1/2}\mathbb{E}_{s\sim\rho}\left[\psi_{\omega}(\phi(s,\pi^{\star}(s)))\right]\right\|_{2}\leq\Lambda$, where $\rho$ is the uniform distribution over ${\mathcal{S}}$. At the same time, for any $\theta_{w}$ we have $\left\|\theta_{w}\right\|_{2}\in\Theta_{B}$ when taking $B=1$, $d>6$ and $\Delta<1/(6d)$.
Next, we will show that $C_{\bm{r}}^{\prime}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\Lambda$. By definition, we have
@@
since $a^{\prime}\equiv a_{3}$ by definition of $\mu_{1}$ and $\psi_{\omega}(\phi(\cdot,a_{3}))\equiv 0$. Then, by Section D.1. of (Zhan et al., 2023), we have $C_{\bm{r}}^{\prime}(\mathcal{G}_{\bm{r}},\pi^{\star},\mu_{1},i)\leq\left\|\Sigma_{\mathcal{D}}^{-1/2}\mathbb{E}_{s\sim\rho}\left[\psi_{\omega}(\phi(s,\pi^{\star}(s)))\right]\right\|_{2}\leq\Lambda$. Therefore, combined with Theorem 3.10 of (Zhu et al., 2023), we finished the proof. ∎
Appendix G Proof of Section 3.2
@@
where $\xi:=\max_{x\in[-R_{\max},R_{\max}]}\left|\frac{\Phi^{\prime}(x)}{\Phi(x)}\right|$, which is also defined in Appendix F. ∎
See 3.4
@@
Here, we used $\mathcal{N}_{\mathcal{G}_{\psi_{\widehat{\omega}}}}(1/NN_{p})=\mathcal{N}_{\mathcal{G}_{\psi^{\star}}}(1/NN_{p})$. Now, we get the similar bound with Equation F.3:
@@
G.1 Why do We Provide Algorithm 6?
Given the inherent complexity of this hierarchical optimization problem, which presents more challenges than standard optimization tasks (Anandalingam and Friesz, 1992), we propose a novel algorithm that circumvents the need for explicit reward function estimation in Algorithm 6. Our approach begins by randomly assigning each human user to a cluster. Subsequently, we reassign random human users to the cluster where the policy most effectively maximizes their empirical DPO loss (Equation E.1). Finally, we refine our solution by optimizing the DPO loss function for the selected human users within each cluster, thereby enhancing the overall policy effectiveness.
Appendix H Proof of Section 4
H.1 Six Pivotal Axioms for Reward Aggregation